# Zero-Shot Interactive Text-to-Image Retrieval via Diffusion-Augmented Representations

### v1

@@ Abstract.
Interactive Text-to-Image Retrieval (I-TIR) has emerged as a transformative user interactive ways for applications in domains such as e-commerce and education. Yet, current methodologies predominantly depend on finetuned Multimodal Large Language Models (MLLMs), which face two critical limitations: (1) Finetuning imposes prohibitive computational overhead and long-term maintenance costs. (2) Finetuning narrows the pretrained knowledge distribution of MLLMs, reducing their adaptability to novel scenarios. These issues are exacerbated by the inherently dynamic nature of real-world I-TIR systems, where queries and image databases evolve in complexity and diversity, often deviating from static training distributions. To overcome these constraints, we propose Diffusion Augmented Retrieval (DAR), a paradigm-shifting framework that bypasses MLLM finetuning entirely. DAR synergizes Large Language Model (LLM)-guided query refinement with Diffusion Model (DM)-based visual synthesis to create contextually enriched intermediate representations. This dual-modality approach deciphers nuanced user intent more holistically, enabling precise alignment between textual queries and visually relevant images. Rigorous evaluations across four benchmarks reveal DAR’s dual strengths: (1) Matches state-of-the-art finetuned I-TIR models on straightforward queries without task-specific training. (2) Scalable Generalization: Surpasses finetuned baselines by 7.61% in Hits@10 (top-10 accuracy) under multi-turn conversational complexity, demonstrating robustness to intricate, distributionally shifted interactions. By eliminating finetuning dependencies and leveraging generative-augmented representations, DAR establishes a new trajectory for efficient, adaptive, and scalable cross-modal retrieval systems.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; July 13–18, 2025; PADOVA, ITALYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Novelty in information retrieval
@@ 1. Introduction
Interactive Text-to-Image Retrieval (I-TIR) seeks to identify relevant images for a user through a turn-by-turn dialogue with a conversational agent. This approach allows users to progressively refine their queries with the agent’s guidance, facilitating the retrieval of highly specific results, even when initial queries are vague or simplistic (Levy et al., 2023), as shown in the top-left of Figure 1. This conversational paradigm is increasingly popular, as it is well-suited to content exploration use cases, such as fashion shopping or searching for long-tail content, where the agent can help the user refine their search (Levy et al., 2023; Wu et al., 2021; Saito et al., 2023; Wan et al., 2024; Yuan and Lam, 2021).
Current state-of-the-art solutions to I-TIR typically rely on finetuning Multimodal Large Language Models (MLLMs) for the retrieval task, aiming to bridge the domain gap between MLLM pretraining and retrieval objectives (Lee et al., 2024; Levy et al., 2023). However, we argue that this finetuning is not always needed. First, the one-to-one mapping between text and images that MLLM fine-tuning aims to establish is neither sufficient nor feasible for complex and diverse quires in I-TIR. Second, by restricting the pretrained distribution, finetuning compromises the broader generalizability that MLLMs acquire from pretraining, causing such models to underperform for I-TIR whenever they encounter dialogues outside the finetuning distribution. This motivates our key research question: How can we enhance the generalizability of I-TIR frameworks without additional training?
To address this research question, we begin by revisiting the motivations for fine-tuning MLLMs and examining its associated drawbacks. MLLM pretraining typically uses large-scale, noisy text-image pairs from internet or crowdsourced data to learn a joint embedding space. However, this approach often leads to sparse alignment, where text and visual representations of the same content may have differing embeddings. In contrast, retrieval tasks require a more exact, one-to-one mapping to maximize text-image similarity scores. Although finetuning on retrieval datasets partially addresses this gap, we argue that achieving such a high degree of alignment is inherently challenging in I-TIR. Textual queries often lack the fine-grained detail needed to describe target images. These limitations make it nearly impossible to establish consistent one-to-one mappings between diverse dialogue queries and target images. Indeed, the finetuned mappings remain confined to the training distribution and therefore lack generalizability.
@@ 2.2. Interactive Text-to-Image Retrieval
Recent studies show the potential of LLMs and MLLMs for I-TIR. For instance, ChatIR (Levy et al., 2023) employs LLMs to simulate dialogues between users and answer bots, compensating for the scarcity of specialized text-to-image datasets tailored to I-TIR. Although this approach opens a promising direction for research, it does not address the unique training hurdles posed by I-TIR—specifically, handling highly diverse dialogue inputs. PlugIR (Lee et al., 2024) further advances I-TIR by improving the diversity of top-k results using $k$-means clustering to group similar candidate images and identify representative exemplars. However, this technique incurs additional computational overhead. Moreover, ChatIR and PlugIR, as well as other I-TIR strategies, rely on fine-tuning on small, curated datasets to achieve good results on a specific benchmark (Zhu et al., 2024; Li et al., 2024; Yoon et al., 2024), limiting their ability to generalize to the broad range of real-world dialogues. Consequently, they often fail on out-of-distribution queries, leading to diminished performance in practical settings, as shown in Section 4.3.
To address these challenges, we propose the DAR framework, which prioritizes zero-shot I-TIR performance. By avoiding dataset-specific fine-tuning altogether, DAR avoids the reduced distribution space introduced by smaller, finetuned datasets, thereby achieving superior generalizability.
@@ 3. DAR for Interactive Text-to-Image Retrieval
  • Diffusion Model (DM): DMs provide valuable prior knowledge about text-to-image mappings—information that un-finetuned MLLMs lack. By generating multiple images based on LLM refined prompts, DMs create multifaceted representations of the user’s intent, thereby bridging the domain gap and eliminating the need for fine-tuning MLLMs.
@@ 4.1. Experimental Settings
We adopt the Stable Diffusion 3 model (SD3) (Esser et al., 2024) as the default diffusion model (DM) in our experiments. Additionally, BLIP-3 (Xue et al., 2024) is employed as the Large Language Model (LLM) for reformulating textual content. Using specially designed prompts, BLIP-3 operates under two distinct reformulation pipelines: one for adapting the dialogue and another for producing aligned prompts for the DM, as described in Section 3.2. We empirically set the weighting factors for textual and visual content to $0.7$ and $0.3$, respectively, for the first two dialogue turns. Starting from turn 3, the weights are both set to $0.5$. This strategy balances the influence of textual vs. visual embeddings as the dialogue becomes more dynamic. In all experiments, we fix the number of generated images per turn at three. Code for all experiments is available at https://anonymous.4open.science/r/Diffusion-Augmented-Retrieval-7EF1/README.md.
@@
  • ZS: The zero-shot (ZS) baseline uses BLIP with its original, publicly available pretrained weights  [^2] Available at: https://github.com/salesforce/BLIP, without any fine-tuning on retrieval datasets. This setup captures the common scenario where researchers or practitioners directly rely on publicly released weights without task-specific training.
@@ 4.2. Zero-Shot I-TIR Performance
We first investigate how our proposed DAR framework performs under zero-shot I-TIR conditions. Specifically, BLIP_zs denotes the baseline where BLIP is used with its original, publicly available pretrained weights—without any fine-tuning on retrieval datasets. Our DAR_zs setup similarly employs pretrianed BLIP as the encoder, thereby reflecting a common scenario in which researchers or practitioners rely solely on publicly released weights.
@@ 4.4. Gains and losses of DAR compared to finetuned models
As analyzed in Section 4.3, FLAN_BLIP2 is the most challenging dataset in our experiments, owing to its particularly complex and dynamic interactive dialogues. Notably, DAR_zs (solid blue line) outperforms the finetuned ChatIR (dash green line) model by 4.22% in Hits@10 at turn 10 (Figure 3.d), supporting our hypothesis that fine-tuning MLLMs on limited I-TIR datasets can undermine their ability to handle out-of-distribution queries. Finetuned models like ChatIR perform best on the VisDial dataset (see Figure 3.a), given their extensive fine-tuning on 123k VisDial training samples. Even in this favorable setting, the worst-case of non-finetuned DAR model, DAR_zs, lags behind ChatIR by only 3.01% in Hits@10, all while saving 100% of the fine-tuning time.
@@ 4.5. Compatibility with finetuned Models
4.6. Does Fine-Tuning on Single-Turn Retrieval Datasets Improve Performance?
Given that single-turn and interactive text-to-image retrieval share some similarities at turn 0, this section investigates whether fine-tuning on a single-turn retrieval dataset—specifically, the widely used MSCOCO dataset (Lin et al., 2014)—enhances performance in interactive retrieval tasks.
As shown in all four subfigures of Figure 3, the benefit of fine-tuning on a single-turn retrieval dataset is relatively limited. While finetuned models such as BLIP_cocoft (dash red lines) and DAR_cocoft (solid red lines) exhibit noticeable performance improvements at earlier turns compared to their zero-shot counterparts (BLIP_zs and DAR_zs), this advantage diminishes as the dialogue progresses. By turn 8, the zero-shot model DAR_zs even surpasses the COCO finetuned model DAR_cocoft in performance, as illustrated in Figure 3.d.
These findings show that conventional finetuning strategies on single-turn datasets are insufficient to address the increasing query diversity and complexity inherent in interactive text-to-image retrieval. This underscores the limitations of fine-tuning-based approaches in handling evolving multi-turn interactions. In contrast, our proposed DAR framework effectively bridges this research gap by enhancing zero-shot performance without requiring additional fine-tuning, making it a more scalable and adaptive solution for real-world retrieval scenarios.
@@ 5.4. Generation Overhead
While DAR eliminates the need for expensive fine-tuning, it introduces additional inference-time overhead due to the reformulation and generation of visual content. However, this trade-off is small compared with the benefits in retrieval performance and adaptability. In our experiments using a single Nvidia RTX 4090 GPU, the image generation process takes approximately 5 seconds, while the query reformulation step incurs only 0.5 seconds of additional processing time.
@@ 6. Conclusion
In this work, we introduce DAR, a novel framework that eliminates the need for fine-tuning multimodal large language models (MLLMs) for Interactive Text-to-Image Retrieval (I-TIR), thereby preserving their generalizability. Extensive experiments show that DAR performs competitively with existing I-TIR models, whether they are fine-tuned or pretrained. Our analysis reveals that fine-tuning MLLMs on limited retrieval datasets compromises their ability to handle out-of-distribution queries. Furthermore, our results illustrate that generating multiple intermediate, multi-faceted representations of user intent enables a many-to-one mapping between text and images, effectively accommodating the diverse and complex queries characteristic of I-TIR. This work highlights the potential of diffusion-augmented retrieval and suggests avenues for future exploration in optimizing efficiency, supporting multimodal queries, and extending to real-world applications.

### v2

# Diffusion Augmented Retrieval: A Training-Free Approach to Interactive Text-to-Image Retrieval

@@ Abstract.
Interactive Text-to-image retrieval (I-TIR) is an important enabler for a wide range of state-of-the-art services in domains such as e-commerce and education. However, current methods rely on finetuned Multimodal Large Language Models (MLLMs), which are costly to train and update, and exhibit poor generalizability. This latter issue is of particular concern, as: 1) finetuning narrows the pretrained distribution of MLLMs, thereby reducing generalizability; and 2) I-TIR introduces increasing query diversity and complexity. As a result, I-TIR solutions are highly likely to encounter queries and images not well represented in any training dataset. To address this, we propose leveraging Diffusion Models (DMs) for text-to-image mapping, to avoid finetuning MLLMs while preserving robust performance on complex queries. Specifically, we introduce Diffusion Augmented Retrieval (DAR), a framework that generates multiple intermediate representations via LLM-based dialogue refinements and DMs, producing a richer depiction of the user’s information needs. This augmented representation facilitates more accurate identification of semantically and visually related images. Extensive experiments on four benchmarks show that for simple queries, DAR achieves results on par with finetuned I-TIR models, yet without incurring their tuning overhead. Moreover, as queries become more complex through additional conversational turns, DAR surpasses finetuned I-TIR models by up to 7.61% in Hits@10 after ten turns, illustrating its improved generalization for more intricate queries.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729950isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Novelty in information retrieval
@@ 1. Introduction
Interactive Text-to-Image Retrieval (I-TIR) seeks to identify relevant images for a user through a turn-by-turn dialogue with a conversational agent. This approach allows users to progressively refine their queries with the agent’s guidance, facilitating the retrieval of highly specific results, even when initial queries are vague or simplistic (Levy et al., 2023), as shown in the top-left of Figure 1. This conversational paradigm is increasingly popular, as it is well-suited to content exploration use cases, such as fashion shopping or searching for long-tail content, where the agent can help the user refine their search (Levy et al., 2023; Wu et al., 2021; Saito et al., 2023; Wan et al., 2024; Yuan and Lam, 2021; Yi et al., 2025).
Current state-of-the-art solutions to I-TIR typically rely on finetuning Multimodal Large Language Models (MLLMs) for the retrieval task (Long et al., 2024b), aiming to bridge the domain gap between MLLM pretraining and retrieval objectives (Lee et al., 2024; Levy et al., 2023; Long et al., 2024c, a, d; Ge et al., 2024, 2021). However, we argue that this finetuning is not always needed. First, the one-to-one mapping between text and images that MLLM finetuning aims to establish is neither sufficient nor feasible for complex and diverse queries in I-TIR. Second, by restricting the pretrained distribution, finetuning compromises the broader generalizability that MLLMs acquire from pretraining, causing such models to underperform for I-TIR whenever they encounter dialogues outside the finetuning distribution. This motivates our key research question: How can we enhance the generalizability of I-TIR frameworks without additional training?
To address this research question, we begin by revisiting the motivations for finetuning MLLMs and examining its associated drawbacks. MLLM pretraining typically uses large-scale, noisy text-image pairs from internet or crowdsourced data to learn a joint embedding space. However, this approach often leads to sparse alignment, where text and visual representations of the same content may have differing embeddings. In contrast, retrieval tasks require a more exact, one-to-one mapping to maximize text-image similarity scores. Although finetuning on retrieval datasets partially addresses this gap, we argue that achieving such a high degree of alignment is inherently challenging in I-TIR. Textual queries often lack the fine-grained detail needed to describe target images. These limitations make it nearly impossible to establish consistent one-to-one mappings between diverse dialogue queries and target images. Indeed, the finetuned mappings remain confined to the training distribution and therefore lack generalizability.
@@ 2.2. Interactive Text-to-Image Retrieval
Recent studies show the potential of LLMs and MLLMs for I-TIR. For instance, ChatIR (Levy et al., 2023) employs LLMs to simulate dialogues between users and answer bots, compensating for the scarcity of specialized text-to-image datasets tailored to I-TIR. Although this approach opens a promising direction for research, it does not address the unique training hurdles posed by I-TIR—specifically, handling highly diverse dialogue inputs. PlugIR (Lee et al., 2024) further advances I-TIR by improving the diversity of top-k results using $k$-means clustering to group similar candidate images and identify representative exemplars. However, this technique incurs additional computational overhead. Moreover, ChatIR and PlugIR, as well as other I-TIR strategies, rely on finetuning on small, curated datasets to achieve good results on a specific benchmark (Zhu et al., 2024; Li et al., 2024; Yoon et al., 2024), limiting their ability to generalize to the broad range of real-world dialogues. Consequently, they often fail on out-of-distribution queries, leading to diminished performance in practical settings, as shown in Section 4.3.
To address these challenges, we propose the DAR framework, which prioritizes zero-shot I-TIR performance. By avoiding dataset-specific finetuning altogether, DAR avoids the reduced distribution space introduced by smaller, finetuned datasets, thereby achieving superior generalizability.
@@ 3. DAR for Interactive Text-to-Image Retrieval
  • Diffusion Model (DM): DMs provide valuable prior knowledge about text-to-image mappings—information that un-finetuned MLLMs lack. By generating multiple images based on LLM refined prompts, DMs create multifaceted representations of the user’s intent, thereby bridging the domain gap and eliminating the need for finetuning MLLMs.
@@ 4.1. Experimental Settings
We adopt the Stable Diffusion 3 model (SD3) (Esser et al., 2024) as the default diffusion model (DM) in our experiments. Additionally, BLIP-3 (Xue et al., 2024) is employed as the Large Language Model (LLM) for reformulating textual content. Using specially designed prompts, BLIP-3 operates under two distinct reformulation pipelines: one for adapting the dialogue and another for producing aligned prompts for the DM, as described in Section 3.2. We empirically set the weighting factors for textual and visual content to $0.7$ and $0.3$, respectively, for the first two dialogue turns. Starting from turn 3, the weights are both set to $0.5$. This strategy balances the influence of textual vs. visual embeddings as the dialogue becomes more dynamic. In all experiments, we fix the number of generated images per turn at three. Code for all experiments is available at https://github.com/longkukuhi/Diffusion-Agumented-Retrieval.
@@
  • ZS: The zero-shot (ZS) baseline uses BLIP with its original, publicly available pretrained weights  [^2] Available at: https://github.com/salesforce/BLIP, without any finetuning on retrieval datasets. This setup captures the common scenario where researchers or practitioners directly rely on publicly released weights without task-specific training.
@@ 4.2. Zero-Shot I-TIR Performance
We first investigate how our proposed DAR framework performs under zero-shot I-TIR conditions. Specifically, BLIP_zs denotes the baseline where BLIP is used with its original, publicly available pretrained weights—without any finetuning on retrieval datasets. Our DAR_zs setup similarly employs pretrained BLIP as the encoder, thereby reflecting a common scenario in which researchers or practitioners rely solely on publicly released weights.
@@ 4.4. Gains and losses of DAR compared to finetuned models
As analyzed in Section 4.3, FLAN_BLIP2 is the most challenging dataset in our experiments, owing to its particularly complex and dynamic interactive dialogues. Notably, DAR_zs (solid blue line) outperforms the finetuned ChatIR (dash green line) model by 4.22% in Hits@10 at turn 10 (Figure 3.d), supporting our hypothesis that finetuning MLLMs on limited I-TIR datasets can undermine their ability to handle out-of-distribution queries. Finetuned models like ChatIR perform best on the VisDial dataset (see Figure 3.a), given their extensive finetuning on 123k VisDial training samples. Even in this favorable setting, the worst-case of non-finetuned DAR model, DAR_zs, lags behind ChatIR by only 3.01% in Hits@10, all while saving 100% of the finetuning time.
@@ 4.5. Compatibility with finetuned Models
4.6. Does finetuning on Single-Turn Retrieval Datasets Improve Performance?
Given that single-turn and interactive text-to-image retrieval share some similarities at turn 0, this section investigates whether finetuning on a single-turn retrieval dataset—specifically, the widely used MSCOCO dataset (Lin et al., 2014)—enhances performance in interactive retrieval tasks.
As shown in all four subfigures of Figure 3, the benefit of finetuning on a single-turn retrieval dataset is relatively limited. While finetuned models such as BLIP_cocoft (dash red lines) and DAR_cocoft (solid red lines) exhibit noticeable performance improvements at earlier turns compared to their zero-shot counterparts (BLIP_zs and DAR_zs), this advantage diminishes as the dialogue progresses. By turn 8, the zero-shot model DAR_zs even surpasses the COCO finetuned model DAR_cocoft in performance, as illustrated in Figure 3.d.
These findings show that conventional finetuning strategies on single-turn datasets are insufficient to address the increasing query diversity and complexity inherent in interactive text-to-image retrieval. This underscores the limitations of finetuning-based approaches in handling evolving multi-turn interactions. In contrast, our proposed DAR framework effectively bridges this research gap by enhancing zero-shot performance without requiring additional finetuning, making it a more scalable and adaptive solution for real-world retrieval scenarios.
@@ 5.4. Generation Overhead
While DAR eliminates the need for expensive finetuning, it introduces additional inference-time overhead due to the reformulation and generation of visual content. However, this trade-off is small compared with the benefits in retrieval performance and adaptability. In our experiments using a single Nvidia RTX 4090 GPU, the image generation process takes approximately 5 seconds, while the query reformulation step incurs only 0.5 seconds of additional processing time.
@@ 6. Conclusion
In this work, we introduce DAR, a novel framework that eliminates the need for finetuning multimodal large language models (MLLMs) for Interactive Text-to-Image Retrieval (I-TIR), thereby preserving their generalizability. Extensive experiments show that DAR performs competitively with existing I-TIR models, whether they are fine-tuned or pretrained. Our analysis reveals that finetuning MLLMs on limited retrieval datasets compromises their ability to handle out-of-distribution queries. Furthermore, our results illustrate that generating multiple intermediate, multi-faceted representations of user intent enables a many-to-one mapping between text and images, effectively accommodating the diverse and complex queries characteristic of I-TIR. This work highlights the potential of diffusion-augmented retrieval and suggests avenues for future exploration in optimizing efficiency, supporting multimodal queries, and extending to real-world applications.